OpenStudy (anonymous):

Can someone please help? Much appreciated <3

OpenStudy (anonymous):

Solve for x and y such that ABCD is a parallelogram.

OpenStudy (tkhunny):

Really Big Hint: Oppostive Angles are Congruent.

OpenStudy (anonymous):

So the answer is 3y-20?

OpenStudy (tkhunny):

No. Can you identify Opposite angles? \(4x = 4y + 4\) You show me the other one.

OpenStudy (anonymous):

2x + 6 = 3y - 20?

OpenStudy (tkhunny):

Perfect. We have: \(4x = 4y+4\) \(2x + 6 = 3y - 20\) Having said that, we can abandon the geometry and work with the algebra problem. Can you solve this system of equations?

OpenStudy (anonymous):

6x + 6 = 7y - 20?

OpenStudy (tkhunny):

I'll take that as a "no". You should have studied several methods to solve such systems. "Substitution" "Matrix Inverse" "Equation Addition" "Determinants" Maybe you haven't heard of all of those. Any sound familiar?

OpenStudy (anonymous):

no haha

OpenStudy (tkhunny):

Well, that's bad, but you did seem to want to add he two equations, so let's figure out how to get that method to do somehting for us. First, rearrange to a consistent form. I'll do the first one. 4x=4y+4 can be written as \(4x - 4y = 4\) You write the other one in a consistent format. x first, then y then = soem constant.